OpenStudy (anonymous):

What is the frequency of the function y = -5cos6x?

OpenStudy (kropot72):

You can use the following equation \[\large y=A \cos Bx\] where the period \[\large T=\frac{2 \pi}{B}\] to find the period of the given function by plugging in values, giving the period \[\large T=\frac{2\pi}{6}\] The frequency is the reciprocal of the period. Therefore \[\large f=\frac{6}{2\pi}\]

OpenStudy (kropot72):

which can be simplified to \[\large f=\frac{3}{\pi}\]

OpenStudy (anonymous):

thank you!!!@kropot72

OpenStudy (anonymous):

no i have a boyfriend @David27

OpenStudy (kropot72):

You're welcome :)